corporation that owns apartment complexes wishes to estimate the average length of time residents remain in the same apartment before moving out. A sample of 150 rental contracts gave a mean length of occupancy of 3.7 years with standard deviation 1.2 years. Construct a 95% confidence interval for the mean length of occupancy of apartments owned by this corporation.

1
Expert's answer
2022-05-05T09:43:01-0400

The critical value for "\alpha = 0.05" and "df = n-1 = 149" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 1.976013."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}} ,\bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(3.7-1.976013\times\dfrac{1.2}{\sqrt{150}},"

"3.7+1.976013\times\dfrac{1.2}{\sqrt{150}})"

"=(3.5064,3.8936)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "3.5064 < \mu < 3.8936," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(3.5064, 3.8936)."